OpenStudy (anonymous):

Help with solving simple rational equations? I have two that I need to do. I got one right but I'm stuck on the other one.

OpenStudy (anonymous):

1) (x + 4)(x - 3) / (x - 3) = 15 x + 4 = 15 = 11 x = 11 2) 2(x2 + 4x - 5) / 2(x2 + 7x + 10) = 4 2(x + 5)(x - 1) / 2(x + 5)(x + 2) = 4 (x - 1)/(x + 2) = 4 x - 1 = 4, x = 5. . . Plug and chug and you get 32! x + 2 = 5, x = 3. . . . Plug and chug and you get 80! I disagree with you. When you got to (x - 1)/(x + 2) = 4, I would've cross multiplied. Would've made a simple rational equation. I handed these into my instructor. He said the first one was good, but he left a comment on the second one. What did he mean? How would I fix it?

OpenStudy (anonymous):

Any help would be greatly appreciated! :-)

OpenStudy (anonymous):

Still stuck on this! haha

OpenStudy (anonymous):

\[\frac{ x-1 }{ x+2 } = 4 ....Cross Multiply.... (x-1) = 4x+8 => 3x = -9 => x = -3\]